# EigenSR: Eigenimage-Bridged Pre-Trained RGB Learners for Single Hyperspectral Image Super-Resolution

### v1

@@ Abstract
Single hyperspectral image super-resolution (single-HSI-SR) aims to improve the resolution of a single input low-resolution HSI. Due to the bottleneck of data scarcity, the development of single-HSI-SR lags far behind that of RGB natural images. In recent years, research on RGB SR has shown that models pre-trained on large-scale benchmark datasets can greatly improve performance on unseen data, which may stand as a remedy for HSI. But how can we transfer the pre-trained RGB model to HSI, to overcome the data-scarcity bottleneck? Because of the significant difference in the channels between the pre-trained RGB model and the HSI, the model cannot focus on the correlation along the spectral dimension, thus limiting its ability to utilize on HSI. Inspired by the HSI spatial-spectral decoupling, we propose a new framework that first fine-tunes the pre-trained model with the spatial components (known as eigenimages), and then infers on unseen HSI using an iterative spectral regularization (ISR) to maintain the spectral correlation. The advantages of our method lie in: 1) we effectively inject the spatial texture processing capabilities of the pre-trained RGB model into HSI while keeping spectral fidelity, 2) learning in the spectral-decorrelated domain can improve the generalizability to spectral-agnostic data, and 3) our inference in the eigenimage domain naturally exploits the spectral low-rank property of HSI, thereby reducing the complexity. This work bridges the gap between pre-trained RGB models and HSI via eigenimages, addressing the issue of limited HSI training data, hence the name EigenSR. Extensive experiments show that EigenSR outperforms the state-of-the-art (SOTA) methods in both spatial and spectral metrics. Our code will be released.
@@ Introduction
Hyperspectral imaging sensors acquire three-dimensional hyperspectral images (HSIs) by capturing a vector with hundreds of contiguous spectral elements from each pixel within a given scene, offering the perception capacity of various types of real-world materials (Keshava and Mustard 2002; Bioucas-Dias et al. 2013). However, physical restrictions in imaging sensors lead to the final product of low spatial resolution HSI (Hong et al. 2021; Yokoya, Grohnfeldt, and Chanussot 2017). To tackle this restriction, hyperspectral image super-resolution (HSI-SR) task (Vivone 2023) has received extensive attention by injecting abundant spatial details to HSI for spatial resolution improvement. Typically, HSI-SR can be categorized into two branches of methods: fusion-based (Yang et al. 2017a; Liu et al. 2022a; Xu et al. 2022; Peng et al. 2021; Xu et al. 2020) and single-image-based methods (Yuan, Zheng, and Lu 2017b; Wan et al. 2022; Liang et al. 2023; Dong et al. 2021; Zhang et al. 2023). In this work, we focus on the single-HSI-SR task.
Similar to the definition of single RGB image SR (Dong et al. 2015; Zhang et al. 2018a), single-HSI-SR aims to recover spatial details with only a single low-resolution (LR) image. With the booming of deep learning and the development of large-scale benchmark datasets (Deng et al. 2009; Agustsson and Timofte 2017), single RGB image SR has reached a level of maturity, exemplified in the studies that leverage pre-training techniques (Chen et al. 2021; Li et al. 2021; Chen et al. 2023). Unfortunately, such an effective strategy cannot be applied to HSI due to the scarcity of data (Liu et al. 2022b; Miao et al. 2022b). Most existing single-HSI-SR methods can only be trained on a small number of indoor and outdoor images, or on low-resolution sub-images of remote sensing data (Jiang et al. 2020; Wang, Ma, and Jiang 2022; Zhang et al. 2023), thus restricting their capabilities. This scarcity of training data has become the major bottleneck in the field of single-HSI-SR.
Given the lack of a large-scale dataset for pre-training an HSI model, a natural question arises: how can we leverage pre-trained RGB models to feed the data-hungry bottleneck of HSI? Directly transferring the RGB model for HSI-SR is difficult, because the channels are not aligned (3 for RGB and much more for HSI) (Yuan, Zheng, and Lu 2017b; Liang et al. 2023). Thus the RGB model cannot handle the correlation along the spectral dimension.
@@
Single RGB image SR has been a long-standing topic of interest. Compared to convolutional neural networks (CNNs) for SR (Dong et al. 2015; Ledig et al. 2017; Wang et al. 2018), Transformer-based architectures (Liang et al. 2021b; Chen et al. 2022, 2023; Hsu, Lee, and Chou 2024) have recently gained prominence in SR due to their ability to extract non-local spatial features and their scalability. Recent studies have found that self-supervised pre-training on large benchmark datasets, such as ImageNet (Deng et al. 2009), followed by fine-tuning on high-quality domain-specific datasets, such as DIV2K (Agustsson and Timofte 2017) and Flickr2K (Timofte et al. 2017), significantly enhances the Transformer’s power to process spatial information (Chen et al. 2021; Li et al. 2021; Chen et al. 2023). However, HSI lacks such large datasets, which hinders the application of pre-training. Therefore, we introduce pre-trained RGB models for HSI to address this data bottleneck.
@@
Single-HSI-SR not only involves processing spatial information but also focuses on the spectral correlation. Researchers used the 3D-CNN (Li, Wang, and Li 2020b), group convolution (Jiang et al. 2020), spectrum and feature context (Wang, Li, and Li 2020), recurrent mechanism (Wang, Ma, and Jiang 2022), and Transformers (Chen, Zhang, and Zhang 2023; Zhang et al. 2023) to learn the spatial-spectral correlations in HSI in an end-to-end manner. Limited by the data, these models may not reach their full potential.
To address this issue, two interesting attempts (Yuan, Zheng, and Lu 2017b; Liang et al. 2023) have found the effectiveness of leveraging RGB images for HSI-SR, inspired by transfer learning. They trained single-channel models on RGB images and inferred on HSI band-by-band to align the channels, followed by a non-negative matrix factorization (NMF) to ensure the spectral fidelity. However, the band-by-band operation in original HSI space and optimization of NMF may result in high time costs. Unlike them, we perform SR in the eigenimage domain, which more effectively and efficiently leverages large-scale pre-trained models.
@@
In this work, we use the pre-trained Image Processing Transformer (IPT) (Chen et al. 2021) as the pre-trained RGB model, which is based on the basic Vision Transformer architecture. The loss function is the L1 loss. For the settings of fine-tuning, please refer to the Experiments section.
@@
Theorem 1 indicates that each channel $\mathbf{E}_{i,:}$ of $\mathbf{E}$ preserves the spatial structure of the original space $\mathbf{Y}$, and that $\mathbf{D}$ acts on each $\mathbf{E}_{i,:}$. This enables us to fine-tune a single-channel model for $\mathbf{E}$, which has already been pre-trained on RGB datasets, to learn the inverse mapping of the same degradation $\mathbf{D}$. The proof of the Theorem 1 is provided in the Appendix.
@@
The high-quality spectral basis $\mathbf{U}_{HR}$ from $\mathbf{Y}_{HR}$ can be used during fine-tuning. However, during inference on unseen data, only the LR HSI is available. We found that the dominant components of $\mathbf{U}$ (e.g., $\mathbf{U}_{:,1}$ and $\mathbf{U}_{:,2}$) at different downsampling rates are not affected seriously, as shown in Figure 3. This is because most spectral information in $\mathbf{Y}_{LR}$ is preserved despite the spatial degradation, even in the extreme case where $\kappa=32$. Therefore, by assuming that the spectral information in $\mathbf{Y}_{LR}$ and $\mathbf{Y}_{HR}$ is close (Yuan, Zheng, and Lu 2017b), we can extract the spectral information from $\mathbf{Y}_{LR}$ to faithfully reconstruct $\mathbf{Y}_{HR}$. We thereby achieve an efficient inference method EigenSR-$\bm{\alpha}$, through:
@@
where $\lambda\in(0,1]$ is a combination weight constant; $\mathbf{Y}_{comb}^{(i)}$ is the combined result starting from $\mathbf{Y}_{LR}$. The detailed explanation for inference with ISR is summarized in Algorithm 1. We denote this version as EigenSR-$\bm{\beta}$. Please note that EigenSR-$\beta$ takes more inference steps, but provides increased spectral regularization. When we set $\lambda$ and the iteration number $N_{it}$ to 1, EigenSR-$\beta$ degenerates to EigenSR-$\alpha$.
@@
We used the peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) to evaluate the spatial quality of SR HSI, and the spectral angle mapper (SAM) for the spectral quality. When testing without reference image, a no-reference quality measurement score (NRQMS) (Yang et al. 2017b) was used to jointly assess the spectral-spatial quality.
@@
We used two NVIDIA 3090 cards to run the algorithms in PyTorch. We adopted the low-rank adaptation (LoRA) (Hu et al. 2021) for parameter efficient fine-tuning, applying LoRA parameters in parallel with all q and v parameters in the pre-trained Transformer (Zhou et al. 2024), and set the low-rank hyperparameter $r$ to 4 as suggested in the original paper. Thus, our trainable parameters consist only of LoRA and the Head and Tail. We fine-tuned on the ARAD_1K-Train dataset for 2000 epochs and on the RESISC45 dataset for 50 epochs, considering the number of training samples. We used the Adam optimizer with a learning rate of 0.001 and the batch size was 64.
@@ Conclusion
This work presents EigenSR for single-HSI-SR, a method that fully leverages the spatial texture processing capabilities of pre-trained RGB models while maintaining spectral correlations, to address the training data bottleneck of HSI. EigenSR is based on the concept of spatial-spectral decoupling via SVD and consists of two stages. First, we fine-tune a single-channel model on the spatial components, i.e., the eigenimages, allowing the RGB model channels to align with HSI and adapt to the eigenimage domain. Then, during inference on unseen data with the fine-tuned model, we decompose the LR HSI, perform channel-wise SR on its eigenimages, and reconstruct the SR HSI using the spectral components, i.e., the spectral basis, to maintain spectral correlations. Furthermore, we propose ISR to fully exploit the LR HSI and enhance spectral fidelity. We only introduce the pre-trained model in the spatial components, allowing spectral correlation to be preserved through the spectral basis, while the SR in the spatial components enhances the model’s generalization ability. Additionally, by utilizing a small number of low-rank eigenimages, we reduce the inference complexity. Extensive experiments demonstrate that our method outperforms state-of-the-art approaches, proving the effectiveness of introducing pre-trained RGB models to address the data scarcity in HSI.

### v2

@@ Abstract
Single hyperspectral image super-resolution (single-HSI-SR) aims to improve the resolution of a single input low-resolution HSI. Due to the bottleneck of data scarcity, the development of single-HSI-SR lags far behind that of RGB natural images. In recent years, research on RGB SR has shown that models pre-trained on large-scale benchmark datasets can greatly improve performance on unseen data, which may stand as a remedy for HSI. But how can we transfer the pre-trained RGB model to HSI, to overcome the data-scarcity bottleneck? Because of the significant difference in the channels between the pre-trained RGB model and the HSI, the model cannot focus on the correlation along the spectral dimension, thus limiting its ability to utilize on HSI. Inspired by the HSI spatial-spectral decoupling, we propose a new framework that first fine-tunes the pre-trained model with the spatial components (known as eigenimages), and then infers on unseen HSI using an iterative spectral regularization (ISR) to maintain the spectral correlation. The advantages of our method lie in: 1) we effectively inject the spatial texture processing capabilities of the pre-trained RGB model into HSI while keeping spectral fidelity, 2) learning in the spectral-decorrelated domain can improve the generalizability to spectral-agnostic data, and 3) our inference in the eigenimage domain naturally exploits the spectral low-rank property of HSI, thereby reducing the complexity. This work bridges the gap between pre-trained RGB models and HSI via eigenimages, addressing the issue of limited HSI training data, hence the name EigenSR. Extensive experiments show that EigenSR outperforms the state-of-the-art (SOTA) methods in both spatial and spectral metrics.
Codehttps://github.com/enter-i-username/EigenSR
@@ Introduction
Hyperspectral imaging sensors acquire three-dimensional hyperspectral images (HSIs) by capturing a vector with hundreds of contiguous spectral elements from each pixel within a given scene, offering the perception capacity of various types of real-world materials (Keshava and Mustard 2002; Bioucas-Dias et al. 2013). However, physical restrictions in imaging sensors lead to the final product of low spatial resolution HSI (Hong et al. 2021; Yokoya, Grohnfeldt, and Chanussot 2017). To tackle this restriction, hyperspectral image super-resolution (HSI-SR) task (Vivone 2023) has received extensive attention by injecting abundant spatial details to HSI for spatial resolution improvement. Typically, HSI-SR can be categorized into two branches of methods: fusion-based (Xu et al. 2022; Peng et al. 2021; Xu et al. 2020) and single-image-based methods (Wan et al. 2022; Liang et al. 2023; Dong et al. 2021; Zhang et al. 2023). In this work, we focus on the single-HSI-SR task.
Similar to the definition of single RGB image SR (Dong et al. 2015; Zhang et al. 2018a), single-HSI-SR aims to recover spatial details with only a single low-resolution (LR) image. With the booming of deep learning (Bai et al. 2024, 2025; Zhou et al. 2024; Nie, Xie, and Sun 2024; Xie et al. 2023; Nie et al. 2023), and the development of large-scale benchmark datasets (Deng et al. 2009; Agustsson and Timofte 2017), single RGB image SR has reached a level of maturity, exemplified in the studies that leverage pre-training techniques (Chen et al. 2021; Li et al. 2021; Chen et al. 2023). Unfortunately, such an effective strategy cannot be applied to HSI due to the scarcity of data (Liu et al. 2022; Miao et al. 2022). Most existing single-HSI-SR methods can only be trained on a small number of indoor and outdoor images, or on low-resolution sub-images of remote sensing data (Jiang et al. 2020; Zhang et al. 2023), thus restricting their capabilities. This scarcity of training data has become the major bottleneck in the field of single-HSI-SR.
Given the lack of a large-scale dataset for pre-training an HSI model, a natural question arises: how can we leverage pre-trained RGB models to feed the data-hungry bottleneck of HSI? Directly transferring the RGB model for HSI-SR is difficult, because the channels are not aligned (3 for RGB and much more for HSI) (Yuan, Zheng, and Lu 2017; Liang et al. 2023). Thus the RGB model cannot handle the correlation along the spectral dimension.
@@
Single RGB image SR has been a long-standing topic of interest. Compared to convolutional neural networks (CNNs) for SR (Dong et al. 2015; Ledig et al. 2017; Wang et al. 2018), Transformer-based architectures (Liang et al. 2021; Chen et al. 2022, 2023; Hsu, Lee, and Chou 2024) have recently gained prominence in SR due to their ability to extract non-local spatial features and their scalability. Recent studies have found that self-supervised pre-training on large benchmark datasets, such as ImageNet (Deng et al. 2009), followed by fine-tuning on high-quality domain-specific datasets, such as DIV2K (Agustsson and Timofte 2017) and Flickr2K (Timofte et al. 2017), significantly enhances the Transformer’s power to process spatial information (Chen et al. 2021; Li et al. 2021; Chen et al. 2023). However, HSI lacks such large datasets, which hinders the application of pre-training. Therefore, we introduce pre-trained RGB models for HSI to address this data bottleneck.
@@
Single-HSI-SR not only involves processing spatial information but also focuses on the spectral correlation. Researchers used the 3D-CNN (Li, Wang, and Li 2020), group convolution (Jiang et al. 2020), spectrum and feature context (Wang, Li, and Li 2020), recurrent mechanism (Wang, Ma, and Jiang 2022), and Transformers (Chen, Zhang, and Zhang 2023; Zhang et al. 2023) to learn the spatial-spectral correlations in HSI in an end-to-end manner. Limited by the data, these models may not reach their full potential.
To address this issue, two interesting attempts (Yuan, Zheng, and Lu 2017; Liang et al. 2023) have found the effectiveness of leveraging RGB images for HSI-SR, inspired by transfer learning. They trained single-channel models on RGB images and inferred on HSI band-by-band to align the channels, followed by a non-negative matrix factorization (NMF) to ensure the spectral fidelity. However, the band-by-band operation in original HSI space and optimization of NMF may result in high time costs. Unlike them, we perform SR in the eigenimage domain, which more effectively and efficiently leverages large-scale pre-trained models.
@@
In this work, we use the pre-trained Image Processing Transformer (IPT) (Chen et al. 2021) as the pre-trained RGB model. The loss function is the L1 loss. For the settings of fine-tuning, please refer to the Experiments section.
@@
Theorem 1 indicates that each channel $\mathbf{E}_{i,:}$ of $\mathbf{E}$ preserves the spatial structure of the original space $\mathbf{Y}$, and that $\mathbf{D}$ acts on each $\mathbf{E}_{i,:}$. This enables us to fine-tune a single-channel model for $\mathbf{E}$, which has already been pre-trained on RGB datasets, to learn the inverse mapping of the same degradation $\mathbf{D}$. The proof of Theorem 1 is given in the Appendix.
@@
The high-quality spectral basis $\mathbf{U}_{HR}$ from $\mathbf{Y}_{HR}$ can be used during fine-tuning. However, during inference on unseen data, only the LR HSI is available. We found that the dominant components of $\mathbf{U}$ (e.g., $\mathbf{U}_{:,1}$ and $\mathbf{U}_{:,2}$) at different downsampling rates are not affected seriously, as shown in Figure 3. This is because most spectral information in $\mathbf{Y}_{LR}$ is preserved despite the spatial degradation, even in the extreme case where $\kappa=32$. Therefore, by assuming that the spectral information in $\mathbf{Y}_{LR}$ and $\mathbf{Y}_{HR}$ is close (Yuan, Zheng, and Lu 2017), we can extract the spectral information from $\mathbf{Y}_{LR}$ to faithfully reconstruct $\mathbf{Y}_{HR}$. We thereby achieve an efficient method EigenSR-$\bm{\alpha}$, through:
@@
where $\lambda\in(0,1]$ is a weight constant; $\mathbf{Y}_{comb}^{(i)}$ is the combined result starting from $\mathbf{Y}_{LR}$. The detailed explanation is in Algorithm 1. We denote this version as EigenSR-$\bm{\beta}$. Note that EigenSR-$\beta$ takes more inference steps, but has increased spectral regularization. When we set $\lambda$ and the iteration number $N_{it}$ to 1, EigenSR-$\beta$ degenerates to EigenSR-$\alpha$.
@@
We used the peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) to evaluate the spatial quality of SR HSI, and the spectral angle mapper (SAM) for the spectral quality. When testing without reference image, a no-reference quality measurement score (NRQMS) (Yang et al. 2017) was used to jointly assess the spectral-spatial quality.
@@
We used two NVIDIA 3090 cards to run the algorithms in PyTorch. We adopted the low-rank adaptation (LoRA) (Hu et al. 2021) for parameter efficient fine-tuning, applying LoRA parameters in parallel with all q and v parameters in the pre-trained Transformer (Zhou et al. 2024b), and set the low-rank hyperparameter $r$ to 4 as suggested in the original paper. Thus, our trainable parameters consist only of LoRA and the Head and Tail. We fine-tuned on the ARAD_1K-Train dataset for 2000 epochs and on the RESISC45 dataset for 50 epochs, considering the number of training samples. We used the Adam optimizer with a learning rate of 0.001 and the batch size was 64.
@@ Conclusion
This work presents EigenSR for single-HSI-SR, a method that fully leverages the spatial capabilities of pre-trained RGB models while maintaining spectral correlations, to address the data bottleneck of HSI. It is based on the concept of spatial-spectral decoupling via SVD and consists of two stages. First, we fine-tune a single-channel model on the spatial components, i.e., the eigenimages, allowing the RGB model channels to align with HSI and adapt to the eigenimage domain. Then, during inference on unseen data, we decompose the LR HSI, perform channel-wise SR on its eigenimages, and reconstruct the SR HSI using the spectral components, i.e., the spectral basis, to maintain spectral correlations. Furthermore, we propose ISR to fully exploit the LR HSI and enhance spectral fidelity. We only introduce the pre-trained model in the spectral-agnostic spatial components, which enhances the model’s generalization ability. Additionally, by utilizing a small number of low-rank eigenimages, we reduce the inference complexity. Extensive experiments demonstrate that our method outperforms SOTA approaches, proving the effectiveness of introducing pre-trained RGB models to address the data scarcity in HSI.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (U2133211, 62301092, 62301093). Xi Su expresses special gratitude to Dr. Shuanghao Bai for providing valuable feedback on the writing of this paper.